OpenStudy (anonymous):

Simplify the expression. 57 – 12 ÷ 4 x 3

OpenStudy (ghostgate):

57 - 12 = 45 and 4 x 3 = 12 45 / 12 = 3.75 Answer = 3.75

OpenStudy (anonymous):

Ah, thanks.

OpenStudy (anonymous):

Can you help me with another?

OpenStudy (ghostgate):

Sure.

OpenStudy (anonymous):

Evaluate. 6*[7+(15-8)]/3 --------------- Which property justifies this statement? r(s + t) = rs + rt A. associative B. distributive C. symmetric D. transitive ---------- Translate the word phrase into a math expression. six less than eight times a number A. 6n - 8 B. 6 - 8n C. 8n - 6 D. 6 < 8n ---------- Which situation could be solved using this expression? 4k + 2 A. Jason read four more than twice the number of books that Marty read. How many books did Marty read? B. Jason read half as many books as Marty read. How many books did Jason and Marty both read? C. Jason read two more than four times the number of books that Marty read. How many books did Jason read? D. Jason read four times as many books as Marty. Marty read two books. How many books did they read together?